OpenStudy (anonymous):

what is the sine value of 2pi/3?? -1/2 1/2 -sqrt3/2 sqrt3/2

OpenStudy (anonymous):

@mathmale @ganeshie8

OpenStudy (anonymous):

@thomas5267

OpenStudy (anonymous):

if you are not going to memorize it , which you should, use a calculator

OpenStudy (thomas5267):

\[ \sin\left(\frac{2\pi}{3}\right)=\sin(120^\circ)=\sin(60^\circ)=\text{draw triangle} \]